OpenStudy (anonymous):

Factor a^6 + 125...My answer below

OpenStudy (anonymous):

\[(a ^{3} + 25)(a ^{3} + 5 )\]

Parth (parthkohli):

\[\rm a^3 + b^3 = (a + b)(a^2 + ab + b^2)\]

Parth (parthkohli):

\[\rm (a^2)^3 + 5^3\]

OpenStudy (anonymous):

Oh, I see. It's already to be put in the sum of two cubes form.

OpenStudy (anonymous):

Thanks!

Parth (parthkohli):

Yes. :D

OpenStudy (anonymous):

I just get confused on the first factor with such a large exponent :P

Parth (parthkohli):

Sorry, I meant,\[a^3 + b^3 = (a + b)(a^2 - ab + b^2) \]

OpenStudy (anonymous):

Yeah, gotta make sure to remember the slight differences between the two similar cube equations for my quiz tomorrow!

Parth (parthkohli):

lol yup! Another fact: That's known as an identity.

OpenStudy (anonymous):

Oh! Thanks for the terminology! :)

Parth (parthkohli):

Heh, no problem. :)